# DiffDA: a diffusion model for weather-scale data assimilation

### 11 Jan 2024 v1

@@ Abstract
The generation of initial conditions via accurate data assimilation is crucial for reliable weather forecasting and climate modeling. We propose the DiffDA as a machine learning based data assimilation method capable of assimilating atmospheric variables using predicted states and sparse observations. We adapt the pretrained GraphCast weather forecast model as a denoising diffusion model.
Our method applies two-phase conditioning: on the predicted state during both training and inference, and on sparse observations during inference only. As a byproduct, this strategy also enables the post-processing of predictions into the future, for which no observations are available.
Through experiments based on a reanalysis dataset, we have verified that our method can produce assimilated global atmospheric data consistent with observations at 0.25degree resolution. The experiments also show that the initial conditions that are generated via our approach can be used for forecast models with a loss of lead time of at most 24 hours when compared to initial conditions of state-of-the-art data assimilation suites. This enables to apply the method to real world applications such as the creation of reanalysis datasets with autoregressive data assimilation.
@@ 1 Introduction
Modern weather forecasts rely heavily on numerical weather predictions. The quality of these predictions depends on the data assimilation process that is used to generate initial conditions on the model grid, based on sparse observations available at different locations. Errors in initial conditions are one of the main sources for errors in weather forecasts [Bonavita et al., 2016]. Additionally, data assimilation is employed in creating reanalysis datasets, such as the ERA5 dataset, which contains reconstructed historical weather variables as grided fields [Hersbach et al., 2020]. Reanalysis datasets play a central role in weather and climate research [ipc, 2023], and are essential for the training of global, machine learned weather forecast models [Pathak et al., 2022, Bi et al., 2023, Lam et al., 2023].
Various data assimilation methods have been developed and employed to address different characteristics of observation data and system dynamics. Among those, variational data assimilation and ensemble Kalman filter are the two most applied families of methods in operational data assimilation [Bannister, 2017]. The variational method is solving an optimization problem minimizing a cost function which measures the discrepancy between forecast simulations of the previous prediction and the real observations. It requires multiple iterations in which linearized observations and evolution functions are evaluated to compute the gradient of the cost function. The linearized observation and evolution function have to be implemented separately. This adds extra overheads because the linearized functions have similar complexity as the original function, and a lot of efforts have to be invested to maintain the consistency of the original code and the linearized code. The ensemble Kalman filter method updates the state estimation according to the covariance matrix calculated from ensemble simulations. Both approaches are computation intensive as one requires multiple optimization iterations and the other requires multiple ensemble simulations.
Data assimilation tools are becoming a bottleneck in the simulation pipeline. While traditional data assimilation methods are sufficiently competent in operational weather forecasts [Bonavita et al., 2016], their high costs restrict broader adoptions, making them tightly coupled with a specific numerical weather forecast model. This restriction becomes more evident since the explosion of ML weather forecast models [Pathak et al., 2022, Lam et al., 2023, Bi et al., 2023]. Those models claim to be a replacement of the traditional models by achieving competitive or even superior accuracy compared to the best operational model while being orders of magnitude faster. Ironically, they cannot independently make forecasts as they are all trained and evaluated on the ERA5 dataset [Hersbach et al., 2020] which is produced by the traditional data assimilation method together with the numerical forecast model. In fact, data assimilation is the last major obstacle to achieve a pure ML weather forecast pipeline.
@@
Figure 1: Diagram of numerical weather forecast pipeline. It consists of data assimilation, forecast and post-processing. Data assimilation produces gridded values from sparse observations and predicted gridded values from previous time steps. Forecast takes in gridded values and produces predictions in gridded values at future time steps. Post-processing improves forecast results such that it is closer to future observations.
@@
In a probabilistic view, data assimilation can be formulated as sampling from a probability distribution of atmosphere states conditioned on observations and predicted states [Law et al., 2015, Evensen et al., 2022]. Capable of solving the conditional sampling problem, denoising diffusion models [Ho et al., 2020] naturally become a tentative choice for data assimilation. Moreover, the blooming community of diffusion models provides an arsenal of techniques for enforcing conditions of different kinds. In particular, conditioning techniques for in-painting [Lugmayr et al., 2022, Song et al., 2020] and super-resolution [Saharia et al., 2022, Chung et al., 2022] are of special interest, because they are similar to conditioning of observations and predicted states respectively. This denoising diffusion model approach has been applied in relatively small scale data assimilation problems [Rozet and Louppe, 2023a, b, Finn et al., 2023, Andry et al., 2023], but none of them is able to assimilate data with a resolution comparable with the ERA5 dataset (0.25 degree horizontal resolution) thus limiting their use with ML forecast models. Similarly, the denoising diffusion techniques have been applied in weather forecasts [Andrychowicz et al., 2023] and post-processing [Mardani et al., 2023, 2024, Li et al., 2023]. In particular, Price et al. developed the GenCast which is capable of doing ensemble forecast at 1 degree resolution.
@@
In this work, we propose a new approach of data assimilation based on the denoising diffusion model with a focus on weather and climate applications. We are able to scale to 0.25degree resolution with 13 vertical levels by utilizing an established ML forecast model GraphCast [Lam et al., 2023] as the backbone of the diffusion model. As we do not have real-world observations available from the ERA5 dataset, we are using grid columns of the ERA5 reanalysis dataset as proxies for observations. During training, the diffusion model is conditioned with background of previous predictions – meaning the state produced by the forecast model using states of forecasts from earlier initial conditions. During inference, we further condition the model with sparse column observations following Repaint [Lugmayr et al., 2022]. In addition, we use a soft mask and interpolated observations to strengthen the conditioning utilizing the continuity of atmosphere variables. The resulting assimilated data can indeed converge to the ground truth data when increasing the number of sampled columns. More importantly, it is able to be used as the inputs of the forecast model where its 48h forecast error also converge to the 48h forecast error with ground truth inputs. We also test the autoregressive data assimilation to generate reanalysis dataset given a time series of observations and an initial predicted field which turns out to be consistent with the ERA5 dataset.
@@
    We demonstrate a novel machine learning (ML) based data assimilation method capable of assimilating high resolution data. The assimilated data is ready for weather forecast applications.
@@
    We are able to create data assimilation cycles combining our method and a machine learning forecast model. The resulting reanalysis dataset is consistent with the ERA5 dataset.
@@
    We build our method using a neural network backbone from a trained machine learning based forecast model. It is easy to upgrade the backbone with any state-of-the-art model due to the flexibility of our method.
  4. 4.
    Our method opens up the possibility of a pure machine learning weather forecast pipeline consisting of data assimilation, forecasting, and post-processing.
@@ 2.1 Problem formulation
The goal of data assimilation is to reconstruct atmospheric variables on a given grid with $n$ grid points $\mathbf{x}_{i}\in\mathbb{R}^{n}$ at physical time step $i$ given $m$ measurements $\mathbf{y}_{i}=f(\mathbf{x}^{*}_{i}),\mathbf{y}_{i}\in\mathbb{R}^{m},\mathbf{x}_{i}^{*}\in\mathbb{R}^{n}$ where $\mathbf{x}^{*}_{i}$ is the ground truth of atmospheric variables on grid points at time step $i$. In addition, estimated values on grid points $\hat{\mathbf{x}}_{i}=\mathcal{F}(\mathbf{x}_{i-1})$ produced by the forecast model $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ are also provided as the inputs. In a probabilistic view, data assimilation samples from a conditional distribution $p(\mathbf{x}_{i}|\hat{\mathbf{x}}_{i},\mathbf{y}_{i})$ which minimizes the discrepancy between $\mathbf{x}_{i}$ and $\mathbf{x}^{*}_{i}$. To simplify the problem, $f$ is limited to a sparse linear observation operator $\mathbf{y}_{i}=f(\mathbf{x}^{*})=\mathbf{Ax}^{*}$ where $\mathbf{A}$ is a sparse matrix with only one nonzero value in each row. In real world cases, this simplification applies to point observations such as temperature, pressure, and wind speed measurements at weather stations and balloons.
@@ 2.2 Denoising diffusion probabilistic model
The denoising diffusion probabilistic model (DDPM) is a generative model capable of sampling from the probabilistic distribution defined by the training data [Ho et al., 2020]. It is trained to approximate the reverse of the diffusion process where noise is added to a state vector $\mathbf{x}^{0}$ during $N$ diffusion steps resulting in an isotropic Gaussian noise vector $\mathbf{x}^{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote the state vector at physical time step $i$ and diffusion step $j$ with $\mathbf{x}^{j}_{i}$. We write $\mathbf{x}^{j}$ whenever the statement is independent of the physical time step. Note that the diffusion step $j$ and the physical time step $i$ are completely independent of each other.
@@
Figure 2: Architecture of the diffusion based data assimilation method. We take advantage of the input and output shape of the pretrained GraphCast model which takes the state of the atmosphere at two time steps as input. In each iteration of the denoising diffusion process in our method, the adapted GraphCast model takes the predicted state and the assimilated state with noise, and further denoises assimilated state. To enforce the observations at inference time, the denoised state is merged with interpolated observations using a soft mask created by blurring the hard mask of the original observations.
@@ 2.3 Conditioning for predicted state
With the conditioning of the predicted state alone, we can already perform backward diffusion steps using the trained model $\mu_{\theta}$ which results in $\mathbf{x}^{0}\sim p(\mathbf{x}^{0}|\hat{\mathbf{x}})$. This means we corrected, or in other words, “post-processed” the forecast result $\hat{\mathbf{x}}$ to be closer to the ground truth state $\mathbf{x}^{*}$.
@@ 2.4 Conditioning for sparse observations
To avoid those issues, we follow inpainting techniques [Lugmayr et al., 2022, Song et al., 2020] to enforce the conditioning of observations at inference time. Let us start with a simple approach first. It creates a hard mask $\mathbf{m}_{h}\in\{0,1\}^{n}$ to indicate which grid columns of the reanalysis state is observed, where a 1 means the associated value in $\mathbf{x}^{*}$ is observed in $\mathbf{y}$ and vice versa. The mask equals the sum of the columns of the observation matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ where $\mathbf{y}=\mathbf{Ax}^{*}$. During the inference of the diffusion model, a state vector with white Gaussian noise is created and gradually denoised by the denoising diffusion model. However, with $\mathbf{y}$ present, we have better knowledge over the observed locations in the state vector. For those locations, values can be produced by a forward diffusion process from the observation data. We can use the mask $\mathbf{m}_{h}$ to treat the two parts separately and combine them in each iteration of the denoising process (Figure 2).
@@
Here, $\odot$ denotes point-wise multiplication used to filter observed and non-observed values. Although this method can guide diffusion result to the observed values, it performs poorly in practice because only the values at the observation locations are forced to the given values while other values remain unchanged as in the unconditional scenario. This is likely related to the encoding-process-decoding architecture commonly used in diffusion models: the encoding and decoding layers employ pooling to downscale and upscale in spatial dimensions. While it helps to condense information and reduce calculations, it also smears out local details. As a result, the added conditioning information is often lost during this process.
We tackle this issue by using a “soft mixing” instead of a hard one. In the soft mixing, we replace the hard mask $\mathbf{m}_{h}$ with a soft mask $\mathbf{m}_{s}$ by applying a Gaussian blur to $\mathbf{m}_{h}$ with standard deviation $\sigma_{G}$, diameter $d$ and re-normalize it to the range of [0, 1]. In this case, the support region of $\mathbf{m}_{s}$ (where its values are larger than 0) is larger than the support region of the observed values. As the atmosphere variables are relatively continuous over space, we interpolate the observed values to fill the support region of $\mathbf{m}_{s}$. While we interpolate all the atmospheric variables in the same way assuming they have similar continuity, it is also possible to adjust the area of support regions for each variable according to its continuity. Thus the inference iteration becomes:
$\begin{split}\mathbf{m}^{\prime}_{s}&=\text{Gaussianblur}(\mathbf{m}_{h},\sigma_{G},d)\\ \mathbf{m}_{s}&=\mathbf{m}^{\prime}_{s}/\max{(\mathbf{m}^{\prime}_{s})}\\ {\mathbf{x}^{*}}^{\prime}&=\text{Interpolate}(\mathbf{m}_{h}\odot\mathbf{x}^{*},d)\\ \mathbf{x}^{j-1}_{\text{known}}&\sim\mathcal{N}(\sqrt{\bar{\alpha}^{j-1}}{\mathbf{x}^{*}}^{\prime},(1-\bar{\alpha}^{j-1})\mathbf{I})\\ \mathbf{x}^{j-1}_{\text{unknown}}&\sim\mathcal{N}(\mu_{\theta}(\mathbf{x}^{j},j),\frac{1-\bar{\alpha}_{j-1}}{1-\bar{\alpha}_{j}}\beta_{t}\mathbf{I})\\ \mathbf{x}^{j-1}&=\mathbf{m}_{s}\odot\mathbf{x}^{j-1}_{\text{known}}+(1-\mathbf{m}_{s})\odot\mathbf{x}^{j-1}_{\text{unknown}}.\end{split}$ (4)
In addition, we also applied the resampling technique from [Lugmayr et al., 2022] to further reduce the inconsistency between the known part and unknown part. For each denoising iteration, the resampling technique repeats the iteration $U$ times by pulling the resulting $\mathbf{x}^{j-1}$ back to $\mathbf{x}^{j}$ with the forward diffusion process (missing) 1 and repeating the denoising step (missing) 4. The overall algorithm of applying denosing diffusion model for data assimilation is presented in Algorithm 1.
@@
0:  predicted state $\hat{\mathbf{x}}$, hard mask of observations $\mathbf{m}_{h}$, observation values at grid points $\mathbf{m}_{h}\odot\mathbf{x}^{*}$ (created from observation vector $\mathbf{y}$, and observation operator $\mathbf{A}$), covariance schedule $\beta_{j},j=1,\cdots,N$, Gaussian blur standard deviation $\sigma_{G}$, Gaussian blur diameter $d$, scaling factor $\mathbf{s}$
@@
  $\mathbf{m}^{\prime}_{s}=\text{Gaussianblur}(\mathbf{m}_{h},\sigma_{G},d)$
  $\mathbf{m}_{s}=\mathbf{m}^{\prime}_{s}/\max{(\mathbf{m}^{\prime}_{s})}$

### 05 Mar 2024 v2

@@ Abstract
The generation of initial conditions via accurate data assimilation is crucial for weather forecasting and climate modeling. We propose DiffDA as a denoising diffusion model capable of assimilating atmospheric variables using predicted states and sparse observations. Acknowledging the similarity between a weather forecast model and a denoising diffusion model dedicated to weather applications, we adapt the pretrained GraphCast neural network as the backbone of the diffusion model. Through experiments based on simulated observations from the ERA5 reanalysis dataset, our method can produce assimilated global atmospheric data consistent with observations at ${}^{\circ}$ ($\approx$30km) resolution globally. This marks the highest resolution achieved by ML data assimilation models. The experiments also show that the initial conditions assimilated from sparse observations (less than $0.77\%$ of gridded data) and 48-hour forecast can be used for forecast models with a loss of lead time of at most 24 hours compared to initial conditions from state-of-the-art data assimilation in ERA5. This enables the application of the method to real-world applications, such as creating reanalysis datasets with autoregressive data assimilation.
Data Assimilation, Denoising Diffusion Model
@@ 1 Introduction
Figure 1: Diagram of numerical weather forecast pipeline. It consists of data assimilation, simulation and post-processing. Data assimilation produces gridded values from sparse observations and predicted gridded values from previous time steps. Simulation takes in gridded values and produces predictions in gridded values at future time steps. Post-processing improves prediction so that it is closer to future observations.
Weather forecast plays an important role in our daily life. Due to uncertainties in the weather forecast process, the predicted weather will inevitably deviate from the actual weather. It is necessary to “pull-back” the predicted weather state via data assimilation, making it consistent with observations from weather stations and satellites. Practically, as Figure 1 shows, the data assimilation process generates initial conditions of weather simulation models from predicted weather state and sparse observations made at different locations. The quality of these weather simulation models depends heavily on the data assimilation since errors in initial conditions are one of the main sources for errors in weather forecasts [Bonavita et al., 2016]. Additionally, data assimilation is employed in creating reanalysis datasets, which contain reconstructed historical weather variables as gridded fields. These reanalysis datasets play a central role in weather and climate research [ipc, 2023, Hersbach et al., 2020], and are essential for the training of ML weather forecast models [Pathak et al., 2022, Bi et al., 2023, Lam et al., 2023].
Figure 2: Architecture of the diffusion-based data assimilation method. We take advantage of the input and output shape of the pretrained GraphCast model, which takes the state of the atmosphere at two time steps as input. In each iteration of the denoising diffusion process, the adapted GraphCast model takes the predicted state and the assimilated state with noise, and further denoises the assimilated state. To enforce the observation values at inference time, The denoised state is merged with interpolated observations using a soft mask created by softbleeding the hard mask derived from the original observations.
Various data assimilation methods have been developed and employed to address different characteristics of observation data and system dynamics. Among those, variational data assimilation and ensemble Kalman filter are the two most applied families of methods in operational data assimilation [Bannister, 2017]. The variational method solves an optimization problem by minimizing a cost function that measures the discrepancy between simulations and observations. It requires multiple iterations in which linearized observations and evolution functions are evaluated to compute the gradient of the cost function. The linearized observation and evolution function have to be implemented separately. This adds extra overheads because the linearized functions have similar complexity as the original function, and a lot of effort has to be invested in maintaining the consistency of the original and linearized code. The ensemble Kalman filter method updates the state estimation according to the covariance matrix calculated from ensemble simulations. Both approaches are computationally intensive as one requires multiple optimization iterations and the other requires multiple ensemble simulations.
Data assimilation tools are becoming a bottleneck in the simulation pipeline. While traditional data assimilation methods are sufficiently competent in operational weather forecasts [Bonavita et al., 2016], their high costs restrict broader adoptions, making them tightly coupled with a specific numerical weather forecast model. This restriction becomes more evident since the explosion of ML weather forecast models [Pathak et al., 2022, Lam et al., 2023, Bi et al., 2023]. Those models claim to be a replacement of the traditional models by achieving competitive or even superior accuracy compared to the best operational model while being orders of magnitude faster. Ironically, they cannot independently make forecasts as they are all trained and evaluated on the ERA5 dataset [Hersbach et al., 2020] which is produced by the traditional data assimilation method together with the numerical forecast model.
@@
In a probabilistic view, data assimilation can be formulated as sampling from a probability distribution of atmosphere states conditioned on observations and predicted states [Law et al., 2015, Evensen et al., 2022]. Capable of solving this conditional sampling problem, denoising diffusion models [Ho et al., 2020] naturally become a tentative choice for data assimilation. Moreover, the blooming community of diffusion models provides an arsenal of techniques for enforcing conditions of different kinds. In particular, conditioning techniques for in-painting [Lugmayr et al., 2022, Song et al., 2020] and super-resolution [Saharia et al., 2022, Chung et al., 2022] are of special interest, because they are similar to conditioning of observations and predicted states respectively. This denoising diffusion model approach has been applied in relatively small-scale data assimilation problems [Rozet and Louppe, 2023a, b, Finn et al., 2023, Andry et al., 2023], but none of them can assimilate data with a resolution comparable with the ERA5 dataset (${}^{\circ}$ horizontal resolution) thus limiting their use with ML forecast models. Similarly, the denoising diffusion techniques have been applied in weather forecasts [Price et al., 2023] and post-processing [Mardani et al., 2023, 2024, Li et al., 2023] with either limited resolution or limited region. While not using diffusion techniques, MetNet3 [Andrychowicz et al., 2023] can make regional weather forecasts from gridded weather data and sparse observations. It combines data assimilation with prediction into one end-to-end process.
@@
In this work, we propose a new approach to data assimilation based on the denoising diffusion model with a focus on weather and climate applications. We are able to scale to ${}^{\circ}$ horizontal resolution with 13 vertical levels by utilizing an established ML forecast model GraphCast [Lam et al., 2023] as the backbone of the diffusion model. As we do not have real-world observations available from the ERA5 dataset, we are using grid columns of the ERA5 reanalysis dataset as proxies for observations. During training, the diffusion model is conditioned with the predicted state – meaning the atmospheric state produced by the forecast model from earlier initial conditions. During inference, we further condition the model with sparse column observations following Repaint [Lugmayr et al., 2022]. In addition, we use a soft mask and interpolated observations to strengthen the conditioning utilizing the continuity of atmospheric variables. The resulting assimilated data can converge to the ground truth field when the number of simulated observation columns increases. More importantly, the assimilated data can be used as the input of the forecast model at a loss of lead time not exceeding 24 hours. We also test the autoregressive data assimilation to generate a reanalysis dataset given a series of observations and an initial predicted field.
@@
    We demonstrate a novel ML data assimilation method capable of assimilating high resolution data. The assimilated data is ready for weather forecast applications.
@@
    We are able to create data assimilation cycles combining our method and an ML forecast model. The resulting reanalysis dataset is consistent with the given observation.
@@
    We build our method using a neural network backbone from a trained ML forecast model. It is easy to upgrade the backbone with any state-of-the-art model due to the flexibility of our method.
@@ 2.1 Problem formulation
The goal of data assimilation is to reconstruct atmospheric variables on a fixed grid with $n$ grid points $\mathbf{x}_{i}\in\mathbb{R}^{n}$ at physical time step $i$ given $m$ measurements $\mathbf{y}_{i}=f(\mathbf{x}^{*}_{i}),\mathbf{y}_{i}\in\mathbb{R}^{m},\mathbf{x}_{i}^{*}\in\mathbb{R}^{n}$ where $\mathbf{x}^{*}_{i}$ is the ground truth of atmospheric variables on grid points at time step $i$. In addition, estimated values on grid points $\hat{\mathbf{x}}_{i}=\mathcal{F}(\mathbf{x}_{i-1})$ produced by the forecast model $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ are also provided as the inputs. In a probabilistic view, data assimilation samples from a conditional distribution $p(\mathbf{x}_{i}|\hat{\mathbf{x}}_{i},\mathbf{y}_{i})$ which minimizes the discrepancy between $\mathbf{x}_{i}$ and $\mathbf{x}^{*}_{i}$. To simplify the problem, $f$ is limited to a sparse linear observation operator $\mathbf{y}_{i}=f(\mathbf{x}^{*})=\mathbf{Ax}^{*}$ where $\mathbf{A}$ is a sparse matrix with only one nonzero value in each row. In real-world cases, this simplification applies to point observations such as temperature, pressure, and wind speed measurements at weather stations and balloons.
@@ 2.2 Denoising diffusion probabilistic model
The denoising diffusion probabilistic model (DDPM) is a generative model capable of sampling from the probabilistic distribution defined by the training data [Ho et al., 2020]. It is trained to approximate the reverse of the diffusion process where noise is added to a state vector $\mathbf{x}^{0}$ during $N$ diffusion steps, resulting in an isotropic Gaussian noise vector $\mathbf{x}^{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote the state vector at physical time step $i$ and diffusion step $j$ with $\mathbf{x}^{j}_{i}$. We write $\mathbf{x}^{j}$ whenever the statement is independent of the physical time step. Note that the diffusion step $j$ and the physical time step $i$ are completely independent of each other.
@@ 2.3 Conditioning for predicted state
With the conditioning of the predicted state alone, we can already perform backward diffusion steps using the trained model $\mu_{\theta}$ which results in $\mathbf{x}^{0}\sim p(\mathbf{x}^{0}|\hat{\mathbf{x}})$. This means we corrected, or in other words, “post-processed” the predictions $\hat{\mathbf{x}}$ to be closer to the ground truth state $\mathbf{x}^{*}$.
@@ 2.4 Conditioning for sparse observations
To avoid those issues, we follow inpainting techniques [Lugmayr et al., 2022, Song et al., 2020] to enforce the conditioning of observations at inference time. Let us start with a simple approach first. It creates a hard mask $\mathbf{m}_{h}\in\{0,1\}^{n}$ to indicate which grid columns of the atmospheric state are observed, where a 1 means the associated value in $\mathbf{x}^{*}$ is observed in $\mathbf{y}$ and vice versa. The mask equals the sum of the columns of the observation matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ where $\mathbf{y}=\mathbf{Ax}^{*}$. During the inference of the diffusion model, a state vector with white Gaussian noise is created and gradually denoised by the denoising diffusion model. However, with $\mathbf{y}$ present, we have better knowledge of the observed locations in the state vector: values can be produced by a forward diffusion process from the observation data. We can use the mask $\mathbf{m}_{h}$ to treat the two parts separately and combine them in each iteration of the denoising process (Figure 2).
@@
Here, $\odot$ denotes point-wise multiplication used to filter observed and non-observed values. Although this method can guide diffusion results to the observed values, it performs poorly in practice because only the values at the observation locations are forced to the given values while other values remain unchanged as in the unconditional scenario. This likely relates to the encoding-process-decoding architecture commonly used in diffusion models: the encoding and decoding layers employ pooling to downscale and upscale in spatial dimensions. While it helps to condense information and reduce calculations, it also smears out local details. As a result, the added conditioning information is often lost during this process.
Figure 3: Creating a soft mask from a hard mask using softbleed. Softbleed performs a ($\max$,$\times$)-convolution over the Gaussian kernel and the hard mask.
We tackle this issue using a “soft mixing” instead of a hard one. In the soft mixing, we replace the hard mask $\mathbf{m}_{h}$ with a soft mask $\mathbf{m}_{s}$ by applying softbleeding to $\mathbf{m}_{h}$ (Figure 3) with standard deviation $\sigma_{G}$ (for Gaussian kernel), and diameter $d$. The softbleed process mimics the Gaussian blurring but replaces the convolution with the ($\max$,$\times$)-convolution to ensure those 1-valued points in the hard mask remain the same in the soft mask. The support region of $\mathbf{m}_{s}$ (where its values are larger than 0) is larger than the support region of the observed values. As the atmospheric variables are relatively continuous over space, we interpolate the observed values to fill the support region of $\mathbf{m}_{s}$. There is abundant flexibility in choosing interpolation algorithms. We use the linear interpolation algorithm provided in scipy package [Virtanen et al., 2020]. In summary, the inference iteration becomes:
$\begin{split}\mathbf{m}_{s}&=\text{Softbleed}(\mathbf{m}_{h},\sigma_{G},d)\\ {\mathbf{x}^{*}}^{\prime}&=\text{Interpolate}(\mathbf{m}_{h}\odot\mathbf{x}^{*},d)\\ \mathbf{x}^{j-1}_{\text{known}}&\sim\mathcal{N}(\sqrt{\bar{\alpha}^{j-1}}{\mathbf{x}^{*}}^{\prime},(1-\bar{\alpha}^{j-1})\mathbf{I})\\ \mathbf{x}^{j-1}_{\text{unknown}}&\sim\mathcal{N}(\mu_{\theta}(\mathbf{x}^{j},j),\frac{1-\bar{\alpha}_{j-1}}{1-\bar{\alpha}_{j}}\beta_{t}\mathbf{I})\\ \mathbf{x}^{j-1}&=\mathbf{m}_{s}\odot\mathbf{x}^{j-1}_{\text{known}}+(1-\mathbf{m}_{s})\odot\mathbf{x}^{j-1}_{\text{unknown}}.\end{split}$ (4)
In addition, we also applied the resampling technique from [Lugmayr et al., 2022] to further reduce the inconsistency between the known part and unknown part. For each denoising iteration, the resampling technique repeats the iteration $U$ times by pulling the resulting $\mathbf{x}^{j-1}$ back to $\mathbf{x}^{j}$ with the forward diffusion process (missing) 1 and repeating the denoising step (missing) 4. The overall algorithm of applying the denoising diffusion model for data assimilation is presented in Algorithm 1.
@@
0:  predicted state $\hat{\mathbf{x}}$, hard mask of observations $\mathbf{m}_{h}$, observation values at grid points $\mathbf{m}_{h}\odot\mathbf{x}^{*}$ (created from observation vector $\mathbf{y}$, and observation operator $\mathbf{A}$), covariance schedule $\beta_{j},j=1,\cdots,N$, Gaussian kernel standard deviation $\sigma_{G}$, Gaussian kernel diameter $d$, scaling factor $\mathbf{s}$
@@
  $\mathbf{m}_{s}=\text{Softbleed}(\mathbf{m}_{h},\sigma_{G},d)$